# Estimating treatment effects with competing intercurrent events in randomized controlled trials

### v1

@@ 3.1 Assumptions
We assume the joint distribution of $\{X_{i},A_{i},T_{i}(1),T_{i}(0),C_{i}(1),C_{i}(0),Y_{i}(1),Y_{i}(0)\}$ for patient $i$ is independently and identically distributed from a super population and will drop the subscript $i$ whenever there is no confusion.
We first assume the following ignorability and overlap assumption for the identification of our causal parameter of interest.
@@
We assume the following conditional independence and overlap conditions for the treatment assignment:
@@
    For some constant $\eta\in(0,0.5)$, $\eta<e(X)<1-\eta$ with probability 1, where $e(X)=\textup{pr}(A=1\mid X)$ denotes the propensity score.
Assumption 1 is guaranteed by the design of an RCT, allowing our results to apply directly to RCTs like our motivating example. However, our formulation is more general as it can also accommodate observational studies. Under Assumption 1, we have the following equations for $\tau$
@@
which are the classic outcome regression and the inverse probability weighting identification formulas when we treat the composite potential outcomes $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ as the “new” outcome variable of interest with $Y^{\textup{cp}}=Y1(T>k)$. However, as discussed in Section 2.2, they do not serve as identification formulas anymore with the presence of the competing treatment-unrelated ICEs because $Y1(T>k)$ is no longer fully observed if treatment-unrelated ICE happened before both $T$ and $k$. Therefore, we need further identification assumptions on the time of treatment-unrelated ICE. Specifically, we assume it is independent of all other potential values, including the potential outcomes, which is equivalent to the censoring at random assumption in the survival analysis literature when we treat treatment-unrelated ICEs as censoring. In addition, we need the censoring probability to be bounded away from 1. We summarize them in the following assumption.
Assumption 2 (Time of treatment-unrelated ICE)
We assume the following conditional independence and positivity conditions for the time of treatment-unrelated ICE.
@@
In our motivating example, Assumption 2 states that the time to treatment-unrelated ICEs, such as withdrawal due to relocation, is independent of both the potential outcome of whether a patient will be a responder at week 52 and the time to an adverse event, conditional on observed baseline covariates. In other words, after accounting for factors such as region, baseline corticosteroid use, and the physician global assessment score, the decision to withdraw due to relocation does not provide additional information about the patients’ eventual response or the time to adverse events. Additionally, it assumes that every patient has a nonzero probability of not experiencing a treatment-unrelated ICE.
@@ 3.2 Two identification formulas
Using the composite variable strategy for the ICEs that are related to the treatment, recall that we define the parameter of interest as:
$\tau=\mu_{1}-\mu_{0}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$
Under Assumptions 1 and 2, we show the parameter of interest $\tau$ is nonparametrically identified and provide two identification formulas in the following theorem.
@@
where $\mu_{a}(X)=E(Y\mid T\wedge C>k,X,A=a)$ is the conditional mean of observed outcome among those with no ICE and $A=a$ and $S_{a}(t\mid X)=\textup{pr}(T>t\mid X,A=a)$ is the conditional survival function of treatment-related ICE time among those with $A=a$. Second,
@@
where $e(X)=\textup{pr}(A=1\mid X)$ is the propensity score and $G_{a}(t\mid X)=\textup{pr}(C>t\mid X,A=a)$ for $a=0,1$ is the survival function of $C$ in the subsample $A=a$.
The first formula (2) expresses $\tau$ using a standard outcome regression approach. It models the conditional mean of the outcome among individuals who have not experienced ICE by $k$, adjusting for the probability of this event occurring. Equation (3) provides an alternative identification using inverse probability weighting. It identifies $\tau$ by averaging observed outcomes only among individuals who have not experienced ICE by time $k$ while correcting for the bias due to differences in the probability of observing a composite outcome. The weights adjust for two key factors: $e(X)$ that accounts for difference in treatment assignment probabilities, and $G_{a}(k\mid X)$ that adjusts for the likelihood of censoring in each treatment group. For instance, units with a lower probability of being observed are up-weighted, while those more likely to remain observed are down-weighted.
Both equations (2) and (3) are valid identification formulas because the nuisance parameters involved are either functions of observed data or identified from observed data. More specifically, nuisance parameters $e(X)$ and $\mu_{a}(X)$ are functions of the observed data. The survival functions of the treatment-related time $S_{a}(t\mid X)$ and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$, although not direct functions of observed data, are both identifiable due to the classic result on survival data with censoring at random (Robins and Rotnitzky,, 1992; Robins and Finkelstein,, 2000). For example, $S_{1}(t\mid X)=\textup{pr}(T>t\mid X,A=1)$ for $t\leq k$ can be estimated using observations in the treatment subgroup from treating $T\wedge C\wedge k$ as the observed event time and $1(C\wedge k\geq T)$ as the event status for the treatment-related ICE, and estimating using a Cox proportional hazard model (Cox,, 1972). We can use a similar procedure to estimate $G_{1}(t\mid X)$ by switching the roles of $T$ and $C$, and use similar procedures to estimate $S_{0}(t\mid X)$ and $G_{0}(t\mid X)$ by using observations in control subgroup.
@@ 3.3 Two basic estimators
Identification formulas (2) and (3) in Theorem 1 motivate two basic estimators based on estimating the corresponding nuisance functions. Let $\hat{S}_{a}(t\mid X)$ and $\hat{G}_{a}(t\mid X)$ denote the fitted models for $t\leq k$ and $a=0,1$, $\hat{\mu}_{a}(X)$ denote the fitted outcome models for $a=0,1$, and $\hat{e}(X)$ denote the fitted propensity score model. Plugging in the fitted values of the nuisance parameters and taking empirical analogs of equations (2) and (3) motivate the following two estimators, respectively,
@@
The outcome regression estimator $\hat{\tau}^{\textup{out}}$ is consistent if the outcome model and the survival function are correctly specified. The weighting estimator $\hat{\tau}^{\textup{ipw}}$ is consistent if the propensity score model and the censoring mechanism are correctly specified.
3.4 Propensity score-augmented weighting estimator
Combining the two identification formulas (2) and (3), we have the following augmented weighting identification formula,
@@ 3.4 Propensity score-augmented weighting estimator
Suppose Assumptions 1 and 2 hold, and assume that $G_{a}(k\mid X)$ is correctly specified for $a=0,1$. $\hat{\tau}^{\textup{aug}}$ is a consistent estimator for $\tau$ if either the propensity score model $e(X)$ is correct, or both the outcome regression $\mu_{a}(X)$ and the survival models $S_{a}(k\mid X)$ are correct for $a=0,1$.
@@
Proposition 1 shows that the consistency of $\hat{\tau}^{\textup{aug}}$ depends on the correct model of $G_{a}(k\mid X)$ for $a=0,1$ but remains robust to misspecification of the propensity score model $e(X)$. To build intuition, Equation (4) can be viewed as a projection of the weighting identification formula (3) onto the nuisance tangent space of the propensity score model $A\mid X$, thus we refer to $\hat{\tau}^{\textup{aug}}$ as a propensity score-augmented estimator. However, since it is not further projected onto the tangent space of the $G_{a}(k\mid X)$ model, the estimator does not retain robustness to misspecification of $G_{a}(k\mid X)$. We will further address the issue in Section 4, where we introduce an estimator that is robust to the misspecification of both models.
@@
$\hat{\tau}^{\textup{aug}}$ dominates the weighting estimator $\hat{\tau}^{\textup{ipw}}$ in terms of robustness. The consistency of $\hat{\tau}^{\textup{aug}}$ is contingent on less restrictive requirements on nuisance parameter estimation in the sense that $\hat{\tau}^{\textup{aug}}$ is consistent whenever $\hat{\tau}^{\textup{ipw}}$ is. Consistency of $\hat{\tau}^{\textup{aug}}$ requires the correct specification of the censoring mechanism $G_{a}(k\mid X)$ for $a=0,1$. Given that $G_{a}(k\mid X)$ is correctly modeled, $\hat{\tau}^{\textup{aug}}$ is doubly robust because it is consistent if either the propensity score model is correct, or the outcome regression and the survival models are both correct. There is no clear dominance between $\hat{\tau}^{\textup{aug}}$ and $\hat{\tau}^{\textup{out}}$ as they require consistent estimation of different sets of nuisance parameters. For all three estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, and $\hat{\tau}^{\textup{aug}}$, we can construct variance estimators using nonparametric bootstrap.
4 Another doubly robust and semiparametrically efficient estimator
@@
The propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$ improves robustness. However, it is not robust to the misspecification of the censoring mechanism, nor does it achieve the semiparametric efficiency bound. In this section, we show that $\hat{\tau}^{\textup{aug}}$ can be further improved. We derive the semiparametric efficient influence function (EIF) for $\tau$ to learn the best asymptotic efficiency a consistent estimator of $\tau$ can achieve and propose an asymptotically efficient and doubly robust estimator based on the EIF.
@@
We first describe the full and observed data structures in our setting and introduce some additional notation. Ideally, we want to observe the full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$. The missingness of the full data comes from two strings. First, for a given treatment arm $a=0,1$, the treatment-unrelated ICE time $C(a)$ censors the full data because $Y^{\textup{cp}}(a)$ is only observable if $C(a)\geq\{T(a)\wedge k\}$. Within each treatment group, we do not observe the full data due to censoring and only observe $(\Delta(a)=1\{C(a)\geq T(a)\wedge k\},\tilde{T}(a)=C(a)\wedge T(a)\wedge k,\Delta(a)Y^{\textup{cp}}(a))$. Second, the treatment assignment generates another level of missingness, because, for each observation, we never simultaneously observe both composite potential outcomes $\{Y^{\textup{cp}}(1),Y^{\textup{cp}}(0)\}$ even without censoring. Therefore, the observed data is $O=(A,\Delta=\Delta(A),\tilde{T}=\tilde{T}(A),\Delta Y^{\textup{cp}}=\Delta(A)Y^{\textup{cp}}(A))$.
4.1 EIF and estimator
@@
and thus, the EIF for $\tau$ is $D_{\tau}(O)=D_{1}(O)-D_{0}(O)$, where $\textnormal{d}M_{G_{a}}(t)=1(C\in\textnormal{d}t,\Delta=0)-1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ with $\Lambda_{a}(t\mid X)$ denoting the conditional cumulative hazard function for the censoring $C$ in the treatment group $A=a$ for $a=0,1$.
The $M_{G_{a}}(t)$ in the EIF is the martingale constructed from the censoring counting process. Intuitively, $1(C\in\textnormal{d}t,\Delta=0)$ is the actual observed increment in the censoring counting process at time $t$, which records whether a censoring event has occurred, while $1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ represents the expected increment in the counting process, given the history up to time $t$. The martingale $M_{G_{a}}(t)$ is essentially capturing the difference between the actual observed events and their expected occurrences. The EIF implies another identification formula for $\tau$ by the property that $E\{D_{\tau}(O)\}=0$. Rearranging terms, we have
@@
where the first two lines are the same as in the propensity score-augmented inverse probability weighting identification formula (4). Intuitively, we further augment (4) by the term (7) to achieve robustness to misspecification of the censoring model. The EIFs in Theorem 2 motivate the following estimators for $\tau$,
@@
More explicitly, taking the first modification term in $\hat{\tau}^{\textup{eif}}$ as an example, because the observed time points are discrete, the estimator of the integration part for observation $i$ is
@@
where $\hat{\lambda}_{1}(t\mid X_{i})$ denotes the estimator of the conditional hazard function of censoring $\lambda_{1}(t\mid X)$. The first term in (8) is a summation of the ratio $-\hat{\lambda}_{1}(t\mid X_{i})/\{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})\}$, where the summation is over all observed event time points before $\tilde{T}_{i}$. The second term is 0 for observations that are not right-censored by the treatment-unrelated ICE, and is $1/\{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})\}$ for observations with a treatment-unrelated ICE happened at time $\tilde{T}_{i}$.
To construct the estimator $\hat{\tau}^{\textup{eif}}$, we need to estimate the following nuisance parameters: the propensity score model $e(X)$, the outcome model $\mu_{a}(X)$ for $a=0,1$, the survival function $S_{a}(t\mid X)$, and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. Nonetheless, the consistency of $\hat{\tau}^{\textup{eif}}$ does not require the correct specification of all four nuisance parameters, as shown in Theorem 3.

### v2

@@ 3.1 Assumptions
We assume that the joint distribution of $\{X_{i},A_{i},T_{i}(1),T_{i}(0),C_{i}(1),C_{i}(0),Y_{i}(1),Y_{i}(0)\}$ for patient $i$ is independently and identically distributed from a superpopulation. For notational simplicity, we omit the subscript $i$ when there is no confusion.
We begin by stating the following assumption on treatment assignment.
@@
We assume the following conditional independence and overlap conditions:
@@
    For some constant $\eta\in(0,0.5)$, $\eta<e(X)<1-\eta$ with probability 1, where $e(X)=\textup{pr}(A=1\mid X)$ denotes the propensity score (Rosenbaum and Rubin,, 1983).
Assumption 1 holds in RCTs by design, allowing our results to apply directly to settings such as our motivating example. However, our formulation is more general and can also accommodate cases where treatment assignment is not completely randomized, such as in stratified randomized experiments and observational studies.
Under Assumption 1, the causal estimand $\tau$ can be expressed as:
@@
which correspond to the standard outcome regression and inverse probability weighting identification formulas, respectively, when treating the composite potential outcome $Y^{\textup{cp}}(a)=Y(a){1}\{T(a)>k\}$ as the “new” outcome of interest. However, as discussed in Section 2.2, these expressions no longer serve as feasible identification formulas in the presence of competing treatment-unrelated ICEs. This is because $Y{1}(T>k)$ is not fully observed if a treatment-unrelated ICE occurs.
To address this challenge, we introduce additional assumptions regarding treatment-unrelated ICEs. Specifically, we assume that the potential time of a treatment-unrelated ICE under treatment $a$ is conditionally independent of the potential outcomes and the potential time to treatment-related ICEs, given baseline covariates. This assumption is analogous to the censoring-at-random assumption in survival analysis when treating treatment-unrelated ICEs as censoring events (Robins et al.,, 1994; Tsiatis,, 2006). Additionally, we require that the probability of no occurrence of treatment-unrelated ICE by time $k$, given observed covariates, is strictly positive. These conditions are summarized in the following assumption.
Assumption 2 (Treatment-unrelated ICE)
We assume the following conditional independence and positivity conditions for the time of treatment-unrelated ICE:
@@
In our motivating example, Assumption 2(a) states that the time to a treatment-unrelated ICE, such as treatment discontinuation due to relocation, is independent of both the potential outcome (i.e., whether a patient would be a responder at week 52) and the time to a treatment-related ICE (e.g., an adverse event), conditional on observed baseline covariates. In other words, after accounting for baseline covariates such as geographic region, baseline corticosteroid use, and the physician’s global assessment score, the decision to discontinue treatment due to relocation is assumed to be unrelated to the patient’s potential outcomes or potential timing of adverse events. Assumption 2(b) requires that every patient has a strictly positive probability of not experiencing a treatment-unrelated ICE before time $k$, given their covariates. This standard positivity condition rules out deterministic treatment-unrelated ICE happening in any subpopulation and is reasonable in our application.
@@ 3.2 Two identification formulas
Using the composite variable strategy to handle treatment-related ICEs, recall that we define the causal estimand as in (1). Under Assumptions 1 and 2, we show that the estimand $\tau$ is nonparametrically identifiable and present two identification formulas in the following theorem.
@@
where for $a=0,1$, $\mu_{a}(X)=E(Y\mid T\wedge C>k,X,A=a)$ is the conditional mean of observed outcome among those with no ICE and $A=a$, and $S_{a}(t\mid X)=\textup{pr}(T>t\mid X,A=a)$ is the conditional survival function of treatment-related ICE in treatment arm $a$. Second,
@@
where $e(X)=\textup{pr}(A=1\mid X)$ is the propensity score and for $a=0,1$, $G_{a}(t\mid X)=\textup{pr}(C>t\mid X,A=a)$ is the conditional survival function of treatment-unrelated ICE in treatment arm $a$.
The first identification formula (2) expresses $\tau$ using a standard outcome regression approach. It decomposes the conditional expectation of the composite outcome given covariates and treatment as $E\{YI(T>k)\mid X,A\}=E(Y\mid T>k,X,A)P(T>k\mid X,A)=E(Y\mid T\wedge C>k,X,A)P(T>k\mid X,A)$. This formulation models the conditional mean outcome $Y$ among individuals who do not have any ICE by time $k$, and weights it by the probability of not experiencing a treatment-related ICE by that time. The second identification formula (3) provides an alternative identification strategy based on inverse probability weighting. It identifies $\tau$ by reweighting observed outcomes among individuals who remain free of ICEs up to time $k$, separately within each treatment arm. The weights adjust for the different probability of observing the composite outcome and consist of two components: $e(X)$, which accounts for differences in treatment assignment probabilities, and $G_{a}(k\mid X)$, which adjusts for the probability of remaining free of treatment-unrelated ICEs up to time $k$ within each treatment group. Effectively, individuals with a lower probability of being observed are up-weighted, while those with a higher probability are down-weighted, ensuring an unbiased estimate of the treatment effect.
Both equations (2) and (3) are valid identification formulas, as the nuisance parameters involved are either functions of observed data or identifiable from observed data. Specifically, the propensity score $e(X)$ and the conditional outcome model $\mu_{a}(X)$ are functions of observed data. Although the survival functions for both types of ICEs, $S_{a}(t\mid X)$ for treatment-related ICEs and $G_{a}(t\mid X)$ for treatment-unrelated ICEs, are not direct functions of the observed data, they are identifiable under standard results in survival analysis with censoring at random (Robins and Rotnitzky,, 1992; Robins and Finkelstein,, 2000; Ebrahimi et al.,, 2003). For example, $S_{1}(t\mid X)=\textup{pr}(T>t\mid X,A=1)$ for $t\leq k$ can be estimated using data from the treatment group by viewing $T\wedge C\wedge k$ as the observed event time and using $1(C\wedge k\geq T)$ as the event indicator for the treatment-related ICE. Estimation can then proceed using parametric or semiparametric methods, such as the Cox proportional hazards model (Cox,, 1972), or more flexible approaches (Wolock et al.,, 2024). A similar approach can be used to estimate $G_{1}(t\mid X)$ by switching the roles of $T$ and $C$. The corresponding functions $S_{0}(t\mid X)$ and $G_{0}(t\mid X)$ can be estimated in the same way using data from the control group.
@@ 3.3 Two basic estimators
The identification formulas (2) and (3) in Theorem 1 motivate two basic estimators, each relying on estimates of the corresponding nuisance functions. Let $\hat{S}_{a}(t\mid X)$ and $\hat{G}_{a}(t\mid X)$ denote the fitted survival models for treatment-related and treatment-unrelated ICEs, respectively, for $t\leq k$ and $a=0,1$. Let $\hat{\mu}_{a}(X)$ be the fitted outcome regression model for each treatment arm, and $\hat{e}(X)$ be the fitted propensity score model. By substituting these fitted values into the identification formulas and taking empirical analogs, we obtain the following two estimators:
@@
The outcome regression estimator $\hat{\tau}^{\textup{out}}$ is consistent if both $\hat{\mu}_{a}(X)$ and $\hat{S}_{a}(k\mid X)$ are consistently estimated. The inverse probability weighting estimator $\hat{\tau}^{\textup{ipw}}$ is consistent if both $\hat{e}(X)$ and $\hat{G}_{a}(k\mid X)$ are consistently estimated.
3.4 An augmented estimator
By combining the two identification formulas (2) and (3), we derive the following augmented identification formula:
@@ 3.4 An augmented estimator
Equation (4) can be interpreted both as a modified form of the outcome regression identification formula (2), and as an augmented form of the weighting identification formula in (3). Under the true model at the population level, the correction terms have mean 00, and thus (4) holds by construction. However, it becomes meaningful in the presence of possible model misspecification, where the augmentation plays a crucial role in improving robustness.
@@
Suppose Assumptions 1 and 2 hold, and assume that $G_{a}(k\mid X)$ is correctly specified for $a=0,1$. $\hat{\tau}^{\textup{aug}}$ is a consistent estimator for $\tau$ if either $e(X)$ is correct, or both $\mu_{a}(X)$ and $S_{a}(k\mid X)$ are correct for $a=0,1$.
@@
Proposition 1 shows that the consistency of $\hat{\tau}^{\textup{aug}}$ depends on correctly specifying $G_{a}(k\mid X)$ for $a=0,1$, while it remains robust to misspecification of the propensity score $e(X)$. To build intuition, from the perspective of semiparametric efficiency theory, Equation (4) can be viewed as a projection of the weighting identification formula (3) onto the nuisance tangent space of the propensity score model $A\mid X$. However, because it is not further projected onto the nuisance tangent space of the $G_{a}(k\mid X)$ model, the resulting estimator does not retain robustness to misspecification of $G_{a}(k\mid X)$. We will further address the issue in Section 4, where we introduce an alternative estimator that is robust to the misspecification of both models.
@@
The augmented estimator $\hat{\tau}^{\textup{aug}}$ dominates the weighting estimator $\hat{\tau}^{\textup{ipw}}$ in terms of robustness. The consistency of $\hat{\tau}^{\textup{aug}}$ is contingent on less restrictive requirements on nuisance parameter estimation in the sense that $\hat{\tau}^{\textup{aug}}$ is consistent whenever $\hat{\tau}^{\textup{ipw}}$ is. Consistency of $\hat{\tau}^{\textup{aug}}$ requires the correct specification of the survival function for treatment-unrelated ICE $G_{a}(k\mid X)$ for $a=0,1$. Given that $G_{a}(k\mid X)$ is correctly modeled, $\hat{\tau}^{\textup{aug}}$ is doubly robust because it is consistent if either the propensity score model is correct, or the outcome regression and the survival model for treatment-related ICEs are both correct. There is no clear dominance between $\hat{\tau}^{\textup{aug}}$ and $\hat{\tau}^{\textup{out}}$ as they require consistent estimation of different sets of nuisance parameters. For all three estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, and $\hat{\tau}^{\textup{aug}}$, we can construct variance estimators using nonparametric bootstrap.
4 A semiparametrically efficient and doubly robust estimator
@@
The augmented weighting estimator $\hat{\tau}^{\textup{aug}}$ improves robustness of $\hat{\tau}^{\textup{ipw}}$. However, it is not robust to the misspecification of the survival function for treatment-unrelated ICE, nor does it achieve the semiparametric efficiency bound. In this section, we show that $\hat{\tau}^{\textup{aug}}$ can be further improved. We derive the semiparametric efficient influence function (EIF) for $\tau$ to learn the best asymptotic efficiency a consistent estimator of $\tau$ can achieve and propose an asymptotically efficient and doubly robust estimator based on the EIF.
@@
We first describe the full and observed data structures in our setting and introduce some additional notation. Ideally, we want to observe the full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$. The missingness of the full data comes from two strings. First, for a given treatment arm $a=0,1$, the treatment-unrelated ICE time $C(a)$ censors the full data because $Y^{\textup{cp}}(a)$ is only observable if $C(a)\geq\{T(a)\wedge k\}$. Within each treatment group, we do not observe the full data due to censoring and only observe
$(\Delta(a)=1\{C(a)\geq T(a)\wedge k\},\tilde{T}(a)=C(a)\wedge T(a)\wedge k,\Delta(a)Y^{\textup{cp}}(a)).$
Second, the treatment assignment generates another level of missingness, because, for each observation, we never simultaneously observe both composite potential outcomes $\{Y^{\textup{cp}}(1),Y^{\textup{cp}}(0)\}$ even without censoring. Therefore, the observed data is
$O=(A,\Delta=\Delta(A),\tilde{T}=\tilde{T}(A),\Delta Y^{\textup{cp}}=\Delta(A)Y^{\textup{cp}}(A)).$
4.1 EIF and EIF-based estimator
@@
and thus, the EIF for $\tau$ is $D_{\tau}(O)=D_{1}(O)-D_{0}(O)$, where $\textnormal{d}M_{G_{a}}(t)=1(C\in\textnormal{d}t,\Delta=0)-1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ with $\Lambda_{a}(t\mid X)$ denoting the conditional cumulative hazard function for the treatment-unrelated ICE $C$ in the treatment group $A=a$ for $a=0,1$.
The $M_{G_{a}}(t)$ in the EIF is the martingale constructed from the censoring counting process. Intuitively, $1(C\in\textnormal{d}t,\Delta=0)$ is the actual observed increment in the censoring counting process at time $t$, which records whether a censoring event has occurred, while $1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ represents the expected increment in the counting process, given the history up to time $t$. The martingale $M_{G_{a}}(t)$ captures the difference between the actual observed events and their expected occurrences. The EIF implies another identification formula for $\tau$ by the property that $E\{D_{\tau}(O)\}=0$. Rearranging terms, we have
@@
where the first two lines are the same as in the augmented weighting identification formula (4). Intuitively, we further augment (4) by the last two terms in (7) to achieve robustness to misspecification of the survival function for treatment-unrelated ICE. These augmentation terms are zero at the population level under true $G_{a}(t\mid X)$. However, their empirical counterparts may deviate significantly from zero, providing diagnostic insight into possible misspecification of $G_{a}(t\mid X)$.
The EIFs in Theorem 2 motivate the following estimator for $\tau$,
@@
To gain insights, consider the first augmentation term in $\hat{\tau}^{\textup{eif}}$. Since the observed time points are discrete, the integral for observation $i$ can be estimated as follows:
@@
where $\hat{\lambda}_{1}(t\mid X_{i})$ denotes the estimated conditional hazard function for the treatment-unrelated ICE at time $t$ given covariates $X_{i}$. In (8), the first term is a sum of the ratio $-\hat{\lambda}_{1}(t\mid X_{i})/\{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})\}$, where the summation is over all observed event time points before $\tilde{T}_{i}$. The second term is 0 for individuals who are not right-censored by the treatment-unrelated ICE, and is $1/\{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})\}$ for observations with a treatment-unrelated ICE occurred at time $\tilde{T}_{i}$.
To construct the estimator $\hat{\tau}^{\textup{eif}}$, we need to estimate the following nuisance parameters: the propensity score model $e(X)$, the outcome model $\mu_{a}(X)$ for $a=0,1$, and the survival functions for treatment-related and treatment-unrelated ICEs $S_{a}(t\mid X)$ and $G_{a}(t\mid X)$, respectively, for $t\leq k$ and $a=0,1$. Importantly, as shown in Theorem 3, the consistency of $\hat{\tau}^{\textup{eif}}$ does not require correct specification of all four nuisance parameters.